# Maxwell’s Demon at Work: Efficient Pruning by Leveraging Saturation of Neurons

### v1

@@ Appendix F Pruning Method Ablation
We validate and justify the heuristic choices made for our pruning method via empirical observation exposed in this section. We used the same setup as before for a ResNet-18 trained on CIFAR-10.
@@ F.1 Regularizer choice
In our empirical analysis, we evaluated the effectiveness of various regularizers on the performance of ResNet-18 networks trained on CIFAR-10, using either Adam or SGDM as optimizers. Our findings (Fig. 6) suggest that focusing regularization exclusively on scale parameters yields a more favorable balance between sparsity and performance with both optimizers. While L2 regularization on scale parameters slightly enhances performance, the scenario changes with SGDM, where Lasso regularization on these parameters outperforms others by a wider margin. Consequently, for the sake of simplicity in our experiments, we have chosen to consistently apply Lasso regularization to scale parameters.
@@ F.2 Dynamic Pruning
F.3 Dead Criterion Relaxation
To measure if a minibatch could be used to measure the death state instead of the entire dataset, we tracked the number of dead neurons during training with both metrics in Fig. 12. We used a minibatch containing 512 inputs from the training dataset for the proxy measurement. We can see that both curves closely track each other. More importantly, they match at the end of the training, indicating that overall the same amount of neurons would be removed when performing the death check over the minibatch. Dynamic pruning was disabled for this experiment.
Figure 12: Instead of validating the death state of neurons against the entire training dataset, it proves sufficient to use a smaller dataset. The curves match throughout training, indicating that roughly the same number of neurons are removed with both strategies. Experiments were performed with ResNet-18 and Lasso($\gamma$) regularization on CIFAR-10 for 3 seeds. Left: With Adam. Right: With SGDM.
@@ F.4 Regularization Schedule
F.5 Added Noise
@@ Appendix G Additional Results
We report additional results in this section, including the pruning performance of DemP when pruning a VGG-16 (Fig. 14) trained with both Adam and SGDM and when pruning a ResNet-50 trained with SGDM (Table 3). Results are similar, with DemP outperforming baselines with Adam at high sparsity. However, the performance of DemP with SGDM decay abruptly passed 95% weight sparsity, which may be due to improper tuning of the method when using a one-cycle learning scheduler to train for a small number of epochs (see Appendix H.2). The results with Leaky ReLU are in Fig. 15.
@@
We included results from Lee et al. (2023) and Evci et al. (2020) in Table 3 to better illustrate the trade-off between structured and unstructured pruning methods. While unstructured methods currently offer more potential to maintain performance at higher parameter sparsity, structured methods offer direct speedup advantages.
Table 3: Comparison between different criteria when pruning a ResNet-50 trained on ImageNet with SGDM around 80% (first line) and 90% (second line) weight sparsity. Similarly to what was observed in smaller models, DemP performs similarly to top baselines, outperforming EarlyCrop here by a small margin. Because structured pruning methods do not have precise control of weight sparsity, we report the closest numbers obtained to these target values. $\pm$ indicates the standard deviation, computed from 3 seeds. The sparsity numbers indicate the removed ratio.
Dense 74.98% ±$\pm 0.08$ 1.0x 1.0x (3.15e18) 1.0x (8.2e9)
Structured SNAP 26.05% ±$\pm 0.17$ 36.90% 81.40% 0.54x 0.33x 0.33x
24.84% ±$\pm 0.11$ 56.07% 90.10% 0.47x 0.25x 0.25x
CroPit-S 25.93% ±$\pm 0.06$ 36.90% 81.40% 0.50x 0.31x 0.31x
25.44% ±$\pm 0.05$ 53.20% 89.80% 0.49x 0.26x 0.26x
EarlySNAP 70.06% ±$\pm 0.16$ 38.73% 79.97% 0.87x 0.57x 0.56x
64.25% ±$\pm 0.10$ 57.23% 89.73% 0.78x 0.44x 0.44x
EarlyCroP-S 71.45% ±$\pm 0.29$ 41.13% 79.77% 0.86x 0.61x 0.61x
67.04% ±$\pm 0.18$ 58.77% 90.07% 0.80x 0.48x 0.48x
DemP-Lasso($\gamma$) 72.21% ±$\pm 0.19$ 53.50% 79.43% 0.88x 0.62x 0.58x
67.76% ±$\pm 0.12$ 68.51% 89.73% 0.69x 0.46x 0.41x
Dense 76.67% - - - - -
Dense* 76.8 ±$\pm 0.09$ 1.0x (3.2e18) 1.0x (8.2e9)
Unstructured Mag 75.53% - 80% - - -
Sal 74.93% - 80% - - -
SET* 72.9%±$\pm 0.39$ 80% - 0.23x 0.23x
69.6%±$\pm 0.23$ 90% - 0.10x 0.10x
RigL (ERK)* 75.10%±$\pm 0.05$ 80% - 0.42x 0.42x
73.00%±$\pm 0.04$ 90% - 0.25x 0.24x
values obtained from Lee et al. (2023)
*values obtained from Evci et al. (2020)
@@ G.1 Comparison with Unstructured Methods
We employ the JaxPruner package (Lee et al., 2023) to illustrate further trade-off of our method against some unstructured methods. Our method is capable of achieving similar performance to unstructured ones for the ResNet-18 experiments (Fig. 16). The comparisons with the unstructured methods use their default configuration from JaxPruner, which was tuned for a ResNet-50. We expect their performance on smaller models to be improved by tuning the pruning distribution, the pruning schedule, and the pruning iterations scheme (Lee et al., 2023). However, for those not interested in expensive tuning, our method becomes an interesting default choice.
@@
Figure 16: Comparison between DemP and unstructured pruning methods from Lee et al. (2023), performed on ResNet-18 trained on CIFAR-10 over 3 seeds. Higher sparsities with DemP are obtained by increasing the peak strength of the added scheduled regularization. Left: With Adam, DemP proves more effective than the unstructured methods left at their base configuration, even if it cannot achieve the same granularity in sparsity by being a structured method. Right: With SGDM, DemP performs similarly.
Appendix H Implementation details
H.1 ResNet-18/ResNet-50
@@
We mostly followed the training procedure of Evci et al. (2020) for the ResNet architectures.
@@
ResNet-18. We train all networks for 250 epochs using a batch size of 128. The learning rate is initially set to 0.005 for Adam, to 0.1 for SGDM, and is thereafter divided by 5 every 77 epochs. While varying regularization is used with our method, it is on top of a constant weight decay (0.0005) used across all methods, including ours. Random crop and random horizontal flips are used for data augmentation.
@@
ResNet-50. We trained the ResNet-50 for 100 epochs, with a batch size of 256 instead of 4096. The initial learning rate is set to 0.005, before being decayed by a factor of 10 at epochs 30, 70, and 90. Label smoothing (0.1) and data augmentation (random resize to either 256 $\times$ 256 or 480 $\times$ 480, before randomly cropping to 224 $\times$ 224. Followed by random horizontal flip and input normalization) are also used. We again use Adam and SGDM, using constant weight decay (0.0001) for both.
H.2 VGG-16
@@
We followed a training procedure similar to Rachwan et al. (2022). We used Adam and SGDM with respectively a learning rate of 0.005 or 0.1 and a batch size of 256, with the One Cycle Learning Rate scheduler (Smith & Topin, 2018). The networks are trained for 80 epochs. CIFAR-10 images are normalized and resized to 64 $\times$ 64 before applying random crop and random horizontal flip for data augmentation.
H.3 Structured Methods
@@
We closely reimplement in JAX (Bradbury et al., 2018) the structured methods from Rachwan et al. (2022), keeping all the hyperparameters specific to every method as is. The training hyperparameters are the same as specified in H.1 and H.2.
H.4 Unstructured Methods
@@
For the unstructured methods, we rely on Lee et al. (2023) implementations, using their method’s configuration for pruning a ResNet-50 for all our experiments. The training hyperparameters are the same as specified in H.1 and H.2.

### v2

@@ Appendix F Pruning Method Ablation
We validate and justify the heuristic choices made for our pruning method, summarized in Algorithm 1, via empirical observation exposed in this section. We used the same setup as before for a ResNet-18 trained on CIFAR-10.
@@ F.1 Regularizer choice
In our empirical analysis, we evaluated the effectiveness of various regularizers on the performance of ResNet-18 networks trained on CIFAR-10, using either Adam or SGDM as optimizers. Our findings (Fig. 5) suggest that focusing regularization exclusively on scale parameters yields a more favourable balance between sparsity and performance with both optimizers. While L2 regularization on scale parameters slightly enhances performance, the scenario changes with SGDM, where Lasso regularization on these parameters outperforms others by a wider margin. Consequently, for the sake of simplicity in our experiments, we have chosen to consistently apply Lasso regularization to scale parameters.
@@ F.2 Dynamic Pruning
F.3 Death Criterion Relaxation
The definition we choose for a dead neuron asks for it to be inactive to the entire dataset. In practice, we found that this criterion could be relaxed and defaulted to using 512 (2048 for ImageNet) examples from the training dataset to measure the death state (measuring across multiple minibatches when necessary). Fig. 12 shows that using this proxy for tracking dead units is sufficient. With this relaxation of the criterion, pruning interventions become highly efficient, incurring at most the computational cost of a few forward propagation batches. To measure if a minibatch could be used to measure the death state instead of the entire dataset, we tracked the number of dead neurons during training with both metrics in Fig. 12. We can see that both curves closely track each other. More importantly, they match at the end of the training, indicating that overall the same amount of neurons would be removed when performing the death check over the minibatch. Dynamic pruning was disabled for this experiment.
Figure 12: Instead of validating the death state of neurons against the entire training dataset, it proves sufficient to use a smaller dataset. The curves match throughout training, indicating that roughly the same number of neurons are removed with both strategies. Experiments were performed with ResNet-18 and Lasso($\gamma$) regularization on CIFAR-10 over three seeds. Left: With Adam. Right: With SGDM.
@@ F.4 Regularization Schedule
F.5 Weight Decay
Our method defaults back to traditional regularization, with a term added directly to the loss, as opposed to the weight decay scheme proposed by DBLP:conf/iclr/LoshchilovH19. By doing so, the adaptive term in optimizers takes into account regularization, and neurons move more quickly toward the inactive region. From a pruning perspective, it achieves a higher sparsity than weight decay for the same regularization strength. Note that we are referring here to the added one-cycle regularization on scale parameters, which on top of potential constant weight decay applied as part of the original training recipe.
F.6 Added Noise
@@
F.7 Activation function
Dead neurons are naturally defined with ReLU activation functions, for which neurons can completely deactivate. However, most activation functions, such as Leaky ReLU (Maas13rectifiernonlinearities), also exhibit a “soft" saturated region. We postulate that neurons firing solely from the saturated region do not contribute much to the predictions and can be considered almost dead. We test this hypothesis by employing our method in a network with Leaky ReLU activations (Fig. 15), removing neurons with only negative activation across a large minibatch. Again, our method can outperform baselines with Adam and offers a similar performance with SGDM.
@@ Appendix G Additional Results
We report additional results in this section, including the pruning performance of DemP when pruning a VGG-16 (Fig. 14) trained with both Adam and SGDM and when pruning a ResNet-50 trained with SGDM (Table 2). Results are similar, with DemP outperforming baselines with Adam at high sparsity. However, the performance of DemP with SGDM decay abruptly passed 95% weight sparsity, which may be due to improper tuning of the method when using a one-cycle learning scheduler to train for a small number of epochs (see Appendix I.2). The results with Leaky ReLU are in Fig. 15.
@@
We included results from jaxpruner and evci2020rigging in Table 2 to better illustrate the tradeoff between structured and unstructured pruning methods. While unstructured methods currently offer more potential to maintain performance at higher parameter sparsity, structured methods offer direct speedup advantages.
@@ G.1 Comparison with Unstructured Methods
We employ the JaxPruner package (jaxpruner) to illustrate further tradeoffs of our method against some unstructured methods. Our method is capable of achieving similar performance to unstructured ones for the ResNet-18 experiments (Fig. 16). The comparisons with the unstructured methods use their default configuration from JaxPruner, which was tuned for a ResNet-50. We expect their performance on smaller models to be improved by tuning the pruning distribution, the pruning schedule, and the pruning iterations scheme (jaxpruner). However, for those not interested in expensive tuning, our method becomes an interesting default choice.
@@
Figure 16: Comparison between DemP and unstructured pruning methods from jaxpruner, performed on ResNet-18 trained on CIFAR-10 over three seeds. Higher sparsities with DemP are obtained by increasing the peak strength of the added scheduled regularization. Left: With Adam, DemP proves more effective than the unstructured methods left at their base configuration, even if it cannot achieve the same granularity in sparsity by being a structured method. Right: With SGDM, DemP performs similarly.
G.2 Transformers Network
We explored the compatibility of DemP with transformer architectures, specifically the Vision Transformer (ViT-B/16) (VisionTransformersDosovitskiyB0WZ21) equipped with GeLU activation functions. Our focus was on pruning the MLP layers within the transformer blocks because these layers account for approximately 65% of the total parameters and contain activation functions, making them directly applicable to DemP.
This investigation led to two significant observations that, to the best of our knowledge, have not been previously reported: (a) even when training the "dense" version, 90% of the neurons within the MLP layers die during training, and (b) in the last MLP block, neurons tend to die quickly before gradually reviving. In other layers, neurons that die remain dead, consistent with our previous observations in other architectures. Those observations are reported in Fig.  17.
To accommodate this behavior, we adapted DemP so that pruning begins after neuron revival occurs in the last MLP layer. The exploratory results, reported in Table 3, are based on $\lambda=0$ (no regularization), but with minimal additional noise applied. Inspired by relustrikebackMirzadehAMMTSRF, we also included results where the GeLU activation functions were replaced with ReLU. Although these results are preliminary, they illustrate the potential of DemP for transformer-based architectures. A promising future direction would be to use DemP in conjunction with the approach proposed by relustrikebackMirzadehAMMTSRF, where additional ReLU activations are incorporated into transformer architectures.
Figure 17: This figure illustrates the evolution of dead neurons during the training process for a 12-layer deep Vision Transformer (ViT-B/16). The rightmost plots depict the behavior in the last layer, where most neurons initially die but subsequently revive, exhibiting a qualitatively different pattern from the other layers. The leftmost and center plots show the typical behavior observed in the other layers, which is more consistent with what has been seen in other architectures.
Appendix H Limitations of our approach
The primary limitation of our work is that, unlike most methods, the target sparsity level cannot be specified within the algorithm. Instead, DemP currently relies on a hyperparameter search over the regularization strength ($\lambda$) to achieve the desired sparsity. However, we are confident that our method can be adapted to this paradigm by continuously increasing the regularization until the desired sparsity is achieved before terminating the process. This strategy is similar to the approach described by Wang2021GrowReg. Additionally, while our analysis demonstrated the impact of batch size and learning rate on final sparsity, for practical reasons, we decided to rely entirely on regularization strength to control the sparsity level. This decision may constrain DemP to suboptimality.
In its current form, DemP removes entire channels from convolutional networks. Adopting a more fine-grained approach, such as removing individual filters, could potentially lead to better performance. Combining DemP with unstructured approaches could also result in a more competitive hybrid method.
Finally, DemP relies on activation functions with saturation regions to identify dead neurons. Consequently, DemP cannot be directly applied to prune layers that are not sandwiched between activation functions, such as the attention layers in transformer blocks. This is reflected in our experiments on ViT-B/16 networks (Section 4.3) where we limited DemP to pruning the fully-connected block within the network. Sparse training methods for Transformers remain underexplored and present unique challenges. A detailed study of DemP in this context, while valuable, is beyond the scope of this paper.
Appendix I Implementation details
I.1 ResNet-18/ResNet-50
@@
We mostly followed the training procedure of evci2020rigging for the ResNet architectures.
@@
ResNet-18. We train all networks for 250 epochs using a batch size of 128. The learning rate is initially set to 0.005 for Adam, to 0.1 for SGDM, and is thereafter divided by 5 every 77 epochs. While varying regularization is used with our method, it is on top of a constant weight decay (0.0005) used across all methods, including ours. Random crop and random horizontal flips are used for data augmentation. The training utilized a Nvidia RTX8000 GPU.
@@
ResNet-50. We trained the ResNet-50 for 100 epochs, with a batch size of 256 instead of 4096. The initial learning rate is set to 0.005, before being decayed by a factor of 10 at epochs 30, 70, and 90. Label smoothing (0.1) and data augmentation (random resize to either 256 $\times$ 256 or 480 $\times$ 480, before randomly cropping to 224 $\times$ 224. Followed by random horizontal flip and input normalization) are also used. We again use Adam and SGDM, using constant weight decay (0.0001) for both. Training utilized a Nvidia A100 GPU.
I.2 VGG-16
@@
We followed a training procedure similar to DBLP:conf/icml/RachwanZCGAG22. We used Adam and SGDM with respectively a learning rate of 0.005 or 0.1 and a batch size of 256, with the One Cycle Learning Rate scheduler (smith2018superconvergence). The networks are trained for 80 epochs. CIFAR-10 images are normalized and resized to 64 $\times$ 64 before applying random crop and random horizontal flip for data augmentation. The training utilized a Nvidia RTX8000 GPU.
I.3 Vision Transformer (ViT-B/16)
We followed the training routine of Lasby2023srigl, that is we added data augmentations following the standard TorchVision (torchvision2016) ViT-B/16 training procedure for ImageNet. These augmentations include random cropping, resizing to 224x224 pixels, random horizontal flips, RandAugment (Cubuk2020), and normalization with typical RGB channel mean and standard deviation values. Additionally, we applied either random mixup (zhang2018mixup) or random cutmix (CutMixYun2019), with alpha parameters of 0.2 and 1.0 respectively.
We sampled 16 mini-batch steps with 256 samples per mini-batch and accumulated gradients before applying the optimizer, resulting in an effective mini-batch size of 4096. The model was trained for 150 epochs using an AdamW (DBLP:conf/iclr/LoshchilovH19) optimizer with weight decay, label smoothing, and $\beta_{1}$, $\beta_{2}$ coefficients of 0.3, 0.11, 0.9, and 0.999, respectively. We used cosine annealing with linear warmup for the learning rate scheduler, starting at 9.9e-5 and warming up to 0.003 by epoch 16. Gradients were clipped to a max L2 norm of 1.0. Uniform sparsity was applied across all layers, with $\Delta T$ set to 100 to update connectivity every 100 mini-batch steps. The training utilized a Nvidia A100 GPU.
I.4 Structured Methods
@@
We closely reimplement in JAX (jax2018github) the structured methods from DBLP:conf/icml/RachwanZCGAG22, keeping all the hyperparameters specific to every method as is. The training hyperparameters are the same as specified in I.1 and I.2.
I.5 Unstructured Methods
@@
For the unstructured methods, we rely on jaxpruner implementations, using their method’s configuration for pruning a ResNet-50 for all our experiments. The training hyperparameters are the same as specified in I.1 and I.2.